# Incentivizing Quality Text Generation via Statistical Contracts

### v1

@@ 1 Introduction
Since the computational requirements of modern LLMs exceed the capacity of common devices, language generation tasks are typically delegated to commercial firms (such as AWS and Microsoft Azure), which operate dedicated computational resources to generate text for a price. To address the tension between LLM performance and computational costs, such firms typically propose multiple service options, with each option offering a different trade-off between model quality and cost [1, 7, 37, 38]. Currently, the most common pricing scheme for such services is pay-per-token, in which users agree in advance to pay a fixed rate for each token of text generated by the system [10].
While simple and intuitive, the pay-per-token pricing scheme creates a misalignment of economic incentives between the firms and their consumers, known in the economic literature as moral hazard: As inference is performed internally and a fixed price is agreed upon in advance, rational firms can strategically increase their profit margin by generating text using a cheaper, lower-quality model. Due to the stochastic nature of language generation, consumers may not be able to reliably determine the quality of the model being used, exposing them to this kind of hazard.
Moral hazard is especially prevalent in cases where the text generation task is intricate, and so evaluation is hard: Consider a scenario where a healthcare provider uses conversational AI to summarize medical notes. As medical diagnosis is an intricate and critical task, the healthcare provider wishes the summaries to be generated using the most advanced language model. Under the pay-per-token pricing scheme, the healthcare provider agrees in advance to pay a fixed amount for each token generated. However, it is not hard to imagine that a firm may attempt to increase profit margins by routing some of the summarization requests to one of its cheaper models, instead of the most advanced one, without taking into account their purpose, and knowing that any lower-quality results would be attributed to the stochastic nature of LLM inference.
@@
In the economic literature, the canonical solution to moral hazard problems is pay-for-performance, or P4P [17]. Instead of paying a fixed price for any outcome, the parties agree in advance on a contract that specifies a differential payment scheme – e.g., agreeing in advance that the firm will receive higher pay when the generated text is considered to be high-quality. When designed correctly, contracts incentivize rational agents to invest more effort, thus providing a way to align incentives. Interaction around contracts is modeled as a principal-agent game, where the principal commits to a payment scheme, and the agent responds by rationally selecting a utility-maximizing action. Within this framework, the principal seeks to design a contract which satisfies some notion of optimality, such as requiring the least amount of expected pay (“min-pay contract”), or the lowest budget (“min-budget contract”).
In this work, we extend the theory of contract design, and use it to design optimal pay-for-performance pricing schemes for delegated text generation. Applying contract design to this setting requires us to overcome the challenges of automated evaluation and cost-robustness. The former arises from the fact that pay-for-performance pricing relies on a measure of performance, and the latter is due to the fact that LLM inference is performed internally by the agent, and therefore the principal possesses uncertainty regarding the agent’s true internal costs.
@@
Finally, we evaluate the empirical performance of cost-robust contracts by analyzing LLM evaluation benchmarks for two families of tasks. In the first experiment, we compare the performance of two-outcome contracts across code generation tasks with varying difficulty; results show that what determines the pricing scheme is the relative success rates of the models, not the task difficulty. In the second experiment, we compute multi-outcome contracts for an intricate conversational task evaluated via LLM-as-a-judge. Numerical results show that the optimal monotone cost-robust pricing scheme has an intuitive 3-level structure: pay nothing if the quality is poor, pay extra if it is exceptional, and pay a fixed baseline otherwise. We show our framework’s flexibility by providing a comprehensive comparison across various contract objectives and monotonicity/threshold constraints.
@@ 1.1 Related work
Our main technical tool is algorithmic contract design (see [5, 20, 14] and subsequent works). We note that many works in this area address distributional robustness, e.g. [8], [14] (which also studies approximation guarantees of simple contracts), and the recent [3] which presents a distributionally-robust contract design approach for delegation of learning – but to our knowledge none address cost-robustness. Connections between contract design and statistics have long been known to exist at a high level (see, e.g., [33]), and were recently explored by [6] in the context of adverse selection, and [32] for two-action min-budget contract. From a technical standpoint, our work is closest to [32], which only proves the statistical connection for the special case of two-action min-budget contracts. Additional related work appears in Appendix A.
@@ 2.1 Quality text generation (agent’s perspective)
The core of our setting is a standard language generation task. Let $V$ be a vocabulary of tokens, and denote the set of all token sequences by $V^{*}$. A text generator $g:V^{*}\to V^{*}$ is a mapping from a textual prompt ${\omega_{0}}$ to a response ${\omega_{R}}$. We assume that prompts are sampled from a distribution ${\omega_{0}}\sim D\in{\Delta\left({V^{*}}\right)}$, and denote by $D_{g}$ the distribution of (prompt, response) pairs where the prompt is sampled from $D$ and the response is generated by generator $g$. Given a prompt and generated response, a quality evaluator is a function $q:V^{*}\times V^{*}\to[m]$ which scores the response on a scale of ${1,\ldots,m}$. We use $F_{g}$ to denote the distribution over scores $[m]$ induced by applying the quality evaluator to a random pair $({\omega_{0}},{\omega_{R}})\sim D_{g}$, and $F_{gj}$ to denote the probability of score $j\in[m]$.
@@
Connection to contract design.
@@
To incentivize high quality text generation, the principal commits in advance to a pay-for-performance contract, which specifies the amount of payment to the agent for generating a response with a certain quality. More formally, given a quality evaluator $q$ with an output scale $1,\ldots,m$, a contract $t:[m]\to\mathbb{R}_{\geq 0}$ is a mapping from the estimated quality to the size of the monetary transfer. Note that transfers are non-negative; this standard restriction is known as limited liability, and it mirrors the fact that when a principal hires an agent to perform a task, money flows in one way only (from principal to agent, and not vice versa). If transfers are increasing with score, we say $t$ is a monotone contract. Monotonicity is not without loss of generality, but is a desirable property as monotone contracts are generally simpler and easier to explain [14].
@@
where $({\omega_{0}},{\omega_{R}})\sim D_{i}$ are the prompt and generated response, $t(q({\omega_{0}},{\omega_{R}}))$ is the payment transferred to the agent based on the quality of response, and ${\alpha}_{i}{\left|{{\omega_{R}}}\right|}$ is the agent’s cost of generating the response. We assume the agent is rational and therefore selects, when facing contract $t$, a pure action $\sigma(t)$ which maximizes their expected profit (also known as the agent’s best response):
@@
As is standard in contract theory, we assume the agent breaks ties among different pure best responses consistently and in a way that agrees with the principal’s preferences. [^3] In our context this means that if action $g_{n}$ is a best response for the agent, then the agent will choose $\sigma(t)$ that plays $g_{n}$ with probability 1 (see Section 2.2).
Figure 1 summarizes the interaction described in this section.
@@ 2.2 Designing the contract (principal’s perspective)
We assume throughout that the principal seeks to obtain text generated from the model $g_{n}$, the most advanced model with the (strictly) highest associated cost $c_{n}>c_{n-1}$. We refer to $g_{n}$ as the target action, i.e. the action which the principal wishes to incentivize. [^4] For high-stake tasks such as summarizing medical information, it makes sense to target the most advanced model. In other cases, assuming a target action serves as a stepping stone towards the final contract design: the principal enumerates over different target actions and ultimately chooses the “best” among the resulting designs. Taking the role of the principal, our goal is now to design the “best” contract $t^{*}$ that incentivizes the agent to generate responses using the target model $g_{n}$. This is formalized by the following optimization problem:
@@
where $\left\lVert{t}\right\rVert$ is a norm of $t$ representing the principal’s economic objective (see below), and $\delta_{g_{n}}$ is a point-mass distribution over text generators, supported by the target generator $g_{n}$. We denote the set of contracts incentivizing action $g_{n}$ by ${\mathcal{T}}(g_{n})={\left\{{t\in\mathbb{R}_{\geq 0}^{m}\mid\sigma(t)=\delta_{g_{n}}}\right\}}$.
@@
In eq. 1, different norms $\left\lVert{t}\right\rVert$ correspond to different possible optimization goals of the principal: For example, a contract is min-pay if it incentivizes the target action using minimum total expected payment $\mathbb{E}_{j\sim F_{n}}{\left[{t(j)}\right]}$ among all contracts in ${\mathcal{T}}(g_{n})$ [14]; In eq. 1, this corresponds to the $\ell_{1}$ norm weighted by the quality distribution of the target action. Similarly, a contract is min-budget if it incentivizes the target action using minimum budget $B_{t}=\max_{j}t(j)$ [32]; In eq. 1, this corresponds to the $\ell_{\infty}$ norm. Additionally, we also consider a natural min-variance objective, which was not previously studied to our knowledge. A min-variance contract minimizes the objective $\mathrm{Var}(t)$, correspoding to a weighted $\ell_{2}$ norm (see Section B.7). We also consider approximately-optimal contracts:
@@ 4 Cost-Robust Contracts
  • The sum-optimal statistical contract $B^{*}_{R}(b)\cdot\psi^{*}_{R}$ is $b$-cost-robust with budget $b/(1-R^{*})$, and has the lowest budget among all $b$-cost-robust contracts;
@@
Table 1 summarizes the contract vs. test equivalences arising from Theorem 1. In the special case of contract settings combining (i) a binary action space ($n=2$), (ii) a tight upper bound ($c_{n}=b$), and (iii) a zero-cost action ($c_{1}=0$), the first half of Theorem 1 recovers a recently-discovered correspondence between hypothesis testing and (non-cost-robust) two-action min-budget contracts [32, Theorem 2]. Theorem 1 is more general since it applies to any number of actions as well as to the standard min-pay objective. Thus, Theorem 1 can also be seen as extending the interpretable format of optimal contracts for binary-action settings beyond two actions.
In Section B.2, we derive our main lemmas, which are used in Section B.4 to prove Theorem 1.
@@ 4.1 Additional properties of optimal cost-robust contracts
In this section, we focus for concreteness on min-budget cost-robust contracts, and establish their approximation guarantees as well as their functional form under assumptions (similar results hold for min-pay cost-robust contracts). First, in analogy to [14], a natural question is what are the approximation guarantees of cost-robust contracts. We show (recall Definition 1):
@@
Proof in Section B.5. As a corollary, combining this result with Theorem 1 shows that statistical contracts are approximately optimal in the global sense: For any contract setting $(F,c)$ with corresponding minimax sum-optimal hypothesis test $\psi^{*}_{R}$, the contract $t=\frac{c_{n}-c_{1}}{1-R^{*}}\psi^{*}_{R}$ is $\eta$-optimal with respect to the budget metric and $\eta=\frac{c_{n}-c_{1}}{c_{n}-c_{n-1}}$.
@@
The first result explains the budget: By the minimax principle, there is a “least favorable distribution” (to use terminology from statistics) or, equivalently, a mixed strategy over the rows of $F$ (to use terminology from game theory) such that no test can achieve for it better risk than the minimax risk $R^{*}$. We show that the budget of the optimal cost-robust contract can be interpreted using this distribution. Formally, let $\left\lVert{{F_{1}}-{F_{2}}}\right\rVert_{\mathrm{TV}}$ be the total variation distance between distributions $F_{1},F_{2}$, then the budget is as follows (see Appendix B.5 for a proof):
@@
The distribution $\sum_{i<n}\lambda_{i}F_{i}$ that maximizes the above expression is the least favorable one. Intuitively, the closer it is to the target distribution $F_{n}$, the larger the budget needed for the agent to distinguish among them and prefer the target action. Finally, we add some standard structure to the distribution matrix $F$ to obtain even simpler contract formats:
@@
Intuitively, if $F$ satisfies MLR, then the higher the outcome $j$, the more likely it is to origin from a more costly distribution $F_{i}$ than from $F_{i^{\prime}}$ (recall that costs $c_{i}$ are increasing in $i$). Consider minimax composite hypothesis tests for ${\left\{{F_{k}}\right\}}_{k=1}^{n-1}$, ${\left\{{F_{n}}\right\}}$; if $F$ does not satisfy MLR, optimal such tests may require randomization (i.e., $\psi_{j}\in(0,1)$ for some outcome $j$) and/or non-monotonicity (i.e. $\psi_{j}>\psi_{j^{\prime}}$ for some pair of outcomes $j<j^{\prime}$). However, if MLR holds for $F$, then nice properties (determinism, monotonicity) hold for minimax tests (e.g., by the Karlin-Rubin theorem [30]), and consequently also for cost-robust contracts (see Appendix B.6 for a proof):
@@
For every contract setting with distribution matrix $F$ that satisfies MLR, and with spread of costs $c_{n}-c_{1}\leq b$, the min-budget $b$-cost-robust contract for $F$ is a monotone threshold contract, which pays full budget to the agent for every outcome $j$ above some threshold $j^{*}$. [^6] For min-pay rather than min-budget, under MLR we get a monotone contract with a single positive payment, which is optimal — see [14, Lemma 7].

### v2

@@ 1 Introduction
Due to the high computational requirements of modern LLMs, language generation tasks are typically outsourced to commercial firms which generate text for a fee. These firms either maintain dedicated infrastructure optimized for inference workloads, or act as intermediaries that facilitate access to such resources. To address the tension between performance and computational costs, such firms typically have multiple service options, each offering a different trade-off between model quality and cost [1, 7, 37, 38]. Currently, the most common pricing scheme for such services is pay-per-token, in which users agree in advance to pay a fixed rate for each token of text generated by the system [10].
While simple and intuitive, the pay-per-token pricing scheme creates a misalignment of economic incentives between the firms and their consumers, known in the economic literature as moral hazard: As inference is performed internally and a fixed price is agreed upon in advance, firms can strategically increase their profit margin by generating text using a cheaper, lower-quality model. Due to the stochastic nature of language generation, consumers may not be able to reliably determine the quality of the model being used, exposing them to this kind of hazard.
Moral hazard is especially prevalent in cases where the text generation task is complex, and so evaluation is hard: Consider a scenario where a healthcare provider hires a firm to use conversational AI for summarizing medical notes. As medical diagnosis is an intricate and critical task, the healthcare provider wishes the medical summaries to be generated using the most advanced language model. Under the pay-per-token pricing scheme, the healthcare provider agrees in advance to pay the firm a fixed amount for each token generated. However, it is not hard to imagine that the firm may attempt to increase profit margins by routing some of the summarization requests to cheaper language models, instead of the most advanced one, without taking into account their purpose, and knowing that any lower-quality results would be attributed to the stochastic nature of LLM inference.
@@
In the economic literature, the canonical solution to moral hazard problems is pay-for-performance, or P4P [17]. Instead of paying a fixed price for any outcome, the parties agree in advance on a contract that specifies a differential payment scheme – for example, agreeing in advance that the firm will receive higher pay when the generated text is considered to be of higher quality. When designed correctly, contracts incentivize rational agents to invest more effort, thus providing a way to align incentives. Interaction around contracts is modeled as a principal-agent game, where the principal commits to a payment scheme, and the agent responds by rationally selecting a utility-maximizing action. Within this framework, the principal seeks to design a contract which satisfies some notion of optimality, such as requiring the least amount of expected pay (“min-pay contract”), or the lowest budget (“min-budget contract”).
In this work, we extend the theory of contract design, and use it to design optimal pay-for-performance pricing schemes for delegated text generation. Applying contract design to this setting requires us to overcome the challenges of automated evaluation and cost uncertainty. The former stems from the need for a scalable measure of performance to support pay-for-performance pricing, while the latter arises from the principal’s uncertainty about the agent’s true internal cost structure, as commercial firms often regard operational costs and implementation details as proprietary information.
@@
Finally, we evaluate the empirical performance of cost-robust contracts by analyzing LLM evaluation benchmarks for two families of tasks. In the first experiment, we compare the performance of two-outcome contracts across code generation tasks with varying difficulty; results show that what determines the pricing scheme is the relative success rates of the models, not the task difficulty. In the second experiment, we compute multi-outcome contracts for an intricate conversational task evaluated via LLM-as-a-judge. Numerical results show that the optimal monotone cost-robust pricing scheme has an intuitive 3-level structure: pay nothing if the quality is poor, pay extra if it is exceptional, and pay a fixed baseline otherwise. We show our framework’s flexibility by providing a comprehensive comparison across various contract objectives and simplicity constraints.
@@ 1.1 Related work
Our main technical tool is algorithmic contract design (see [5, 20, 14] and subsequent works). Many works in this area address distributional robustness, e.g. [8], [14] which also studies approximation guarantees of simple contracts, and the recent [3] which presents a distributionally-robust contract design approach for delegation of learning. However, to our knowledge, none address cost-robustness. Connections between contract design and statistics have long been known to exist at a high level (see, e.g., [33]), and were recently explored by [6] in the context of adverse selection, and [32] for two-action min-budget contract. From a technical standpoint, our work is closest to [32], which only proves the statistical connection for the special case of two-action min-budget contracts. Finally, we note that our cost-robustness framework is general, and our characterization results may be of independent interest. Additional related work appears in Appendix A.
@@ 2.1 Quality text generation (agent’s perspective)
The core of our setting is a standard language generation task. Let $V$ be a vocabulary of tokens, and denote the set of all token sequences by $V^{*}$. A text generator $g:V^{*}\to V^{*}$ is a mapping from a textual prompt ${\omega_{0}}$ to a response ${\omega_{R}}$. We assume that prompts are sampled from a distribution ${\omega_{0}}\sim D\in{\Delta\left({V^{*}}\right)}$, and denote by $D_{g}$ the distribution of prompt-response pairs, where the prompt is sampled from $D$ and the response is generated by generator $g$. Given a prompt and generated response, a quality evaluator is a function $q:V^{*}\times V^{*}\to[m]$ which scores the response on a scale of ${1,\ldots,m}$. We use $F_{g}$ to denote the distribution over scores $[m]$ induced by applying the quality evaluator to a random pair $({\omega_{0}},{\omega_{R}})\sim D_{g}$, and $F_{gj}$ to denote the probability of score $j\in[m]$.
@@
As an abstract contract design problem.
@@
To incentivize high quality text generation, the principal commits in advance to a pay-for-performance contract, which specifies the amount of payment to the agent for generating a response with a certain quality. More formally, given a quality evaluator $q$ with an output scale $1,\ldots,m$, a contract $t:[m]\to\mathbb{R}_{\geq 0}$ is a mapping from the estimated quality to the size of monetary transfer. Note that transfers are non-negative; this standard restriction is known as limited liability, and it mirrors the fact that when a principal hires an agent to perform a task, money flows in one way only (from principal to agent, and not vice versa). If transfers are increasing with score, we say $t$ is a monotone contract. Monotonicity is not without loss of generality, but is a desirable property as monotone contracts are generally simpler and easier to explain [14].
@@
where $({\omega_{0}},{\omega_{R}})\sim D_{i}$ are the prompt and generated response, $t(q({\omega_{0}},{\omega_{R}}))$ is the payment transferred to the agent based on the quality of response, and ${\alpha}_{i}{\left|{{\omega_{R}}}\right|}$ is the agent’s cost of generating the response. We assume the agent is rational and therefore selects, when facing contract $t$, an action $\sigma(t)$ which maximizes their expected profit (also known as the agent’s best response):
@@
As is standard in contract theory, we assume the agent breaks ties consistently and in a way that agrees with the principal’s preferences. [^3] In our context, this means that if action $g_{n}$ is a best response for the agent, then the agent will choose $\sigma(t)$ that plays $g_{n}$ with probability 1 (see Section 2.2). The interaction model is summarized in Figure 1.
@@ 2.2 Designing the contract (principal’s perspective)
We assume that the principal seeks to obtain text generated by the model $g_{n}\in{\mathcal{G}}$, the most advanced model with the (strictly) highest associated cost $c_{n}>c_{n-1}$. We refer to $g_{n}$ as the target action, i.e. the action which the principal wishes to incentivize. Taking the role of the principal, our goal is to design the “best” contract $t^{*}$ that incentivizes the agent to generate responses using the target model $g_{n}$. This is formalized by the following optimization problem:
@@
where $\left\lVert{t}\right\rVert$ is a norm of $t$ representing the principal’s economic objective (see below), and $\delta_{g_{n}}$ is a point-mass distribution over text generators, supported by the target generator $g_{n}$. We denote the set of contracts incentivizing action $g_{n}$ by ${\mathcal{T}}(g_{n})={\left\{{t\in\mathbb{R}_{\geq 0}^{m}\mid\sigma(t)=\delta_{g_{n}}}\right\}}$, and further note that the assumption of a single target action serves as a foundational step towards more complex contract design scenarios (see Section B.1).
@@
In eq. 1, different norms $\left\lVert{t}\right\rVert$ correspond to different possible optimization goals of the principal: For example, a contract is min-pay if it incentivizes the target action using minimum total expected payment $\mathbb{E}_{j\sim F_{n}}{\left[{t(j)}\right]}$ among all contracts in ${\mathcal{T}}(g_{n})$ [14]; In eq. 1, this corresponds to the $\ell_{1}$ norm weighted by the quality distribution of the target action. Similarly, a contract is min-budget if it incentivizes the target action using minimum budget $B_{t}=\max_{j}t(j)$ [32]; In eq. 1, this corresponds to the $\ell_{\infty}$ norm. Additionally, we also consider a natural min-variance objective, which was not previously studied to our knowledge. A min-variance contract minimizes the objective $\mathrm{Var}(t)$, corresponding to a weighted $\ell_{2}$ norm. Optimal contracts for these objectives can be computed in polynomial time by solving a corresponding linear or convex-quadratic program (see Section D.1). We also consider approximately-optimal contracts:
@@ 4 Cost-Robust Contracts
  • The sum-optimal statistical contract $B^{*}_{R}(b)\cdot\psi^{*}_{R}$ is $b$-cost-robust with budget $b/(1-R^{*})$, and has the lowest budget among all $b$-cost-robust contracts.
@@
Table 1 summarizes the contract vs. test equivalences arising from Theorem 1. In the special case of contract design settings combining (i) a binary action space ($n=2$), (ii) a zero-cost action ($c_{1}=0$), and (iii) a tight upper bound ($b=c_{2}$), the first half of Theorem 1 recovers a recently-discovered correspondence between hypothesis testing and (non-cost-robust) two-action min-budget contracts [32, Theorem 2]. Theorem 1 is more general since it applies to any number of actions as well as to the standard min-pay objective. Thus, Theorem 1 can also be seen as extending the interpretable format of optimal contracts for binary-action settings beyond two actions.
In Section D.2, we derive our main lemmas, which are used in Section D.4 to prove Theorem 1. Our proofs rely on two main assumptions. The first assumption is that the cost uncertainty bounds are known; as these bounds become looser, the required budget and expected payout increase linearly. The second assumption is that the contract design problem is implementable – i.e., that there exists some contract incentivizing the target action. In particular, implementability holds when text generated by the target model has the highest quality in expectation (see Section C.1). Generally, a contract design problem is implementable when the observed quality distribution of the target model can’t be emulated by a combination of alternative models at a lower cost (see, e.g., [14]).
@@ 4.1 Additional properties of optimal cost-robust contracts
In this section, we focus for concreteness on min-budget cost-robust contracts, and establish their approximation guarantees as well as their functional form under structural assumptions. First, in analogy to [14], it is natural to examine the approximation guarantees of cost-robust contracts. We show (recall Definition 1):
@@
Proof in Section D.5. As a corollary, combining this result with Theorem 1 shows that statistical contracts are approximately optimal in the global sense: For any contract setting $(F,c)$ with corresponding minimax sum-optimal hypothesis test $\psi^{*}_{R}$, the contract $t=\frac{c_{n}-c_{1}}{1-R^{*}}\psi^{*}_{R}$ is $\eta$-optimal with respect to the budget metric and $\eta=\frac{c_{n}-c_{1}}{c_{n}-c_{n-1}}$. We also note that similar results hold for min-pay cost-robust contracts.
@@
The first result explains the budget: By the minimax principle, there is a “least favorable distribution” (to use terminology from statistics) or, equivalently, a mixed strategy over the rows of $F$ (to use terminology from game theory) such that no test can achieve for it better risk than the minimax risk $R^{*}$. We show that the budget of the optimal cost-robust contract can be interpreted using this distribution. Formally, let $\left\lVert{{F_{1}}-{F_{2}}}\right\rVert_{\mathrm{TV}}$ be the total variation distance between distributions $F_{1},F_{2}$, then the budget is as follows (see Appendix D.5 for a proof):
@@
The distribution $\sum_{i<n}\lambda_{i}F_{i}$ that maximizes the above expression is the least favorable one. Intuitively, the closer it is to the target distribution $F_{n}$, the larger the budget needed for the agent to distinguish among them and prefer the target action. We also note that the required budget approaches infinity when the worst-case distribution distance approaches zero, coinciding with the implementability characterization known from the literature (see Section C.1). The second set of results adds standard structure to the distribution matrix $F$ to obtain even simpler contract formats:
@@
Intuitively, if $F$ satisfies MLR, then the higher the outcome $j$, the more likely it is to origin from a more costly distribution $F_{i}$ than from $F_{i^{\prime}}$ (recall that costs $c_{i}$ are increasing in $i$). Consider minimax composite hypothesis tests for ${\left\{{F_{k}}\right\}}_{k=1}^{n-1}$, ${\left\{{F_{n}}\right\}}$; if $F$ does not satisfy MLR, optimal such tests may require randomization (i.e., $\psi_{j}\in(0,1)$ for some outcome $j$) and/or non-monotonicity (i.e. $\psi_{j}>\psi_{j^{\prime}}$ for some pair of outcomes $j<j^{\prime}$). However, if MLR holds for $F$, then nice properties (determinism, monotonicity) hold for minimax tests (e.g., by the Karlin-Rubin theorem [30]), and consequently also for cost-robust contracts (see Appendix D.6 for a proof):
@@
For every contract setting with distribution matrix $F$ that satisfies MLR, and with spread of costs $c_{n}-c_{1}\leq b$, the min-budget $b$-cost-robust contract for $F$ is a monotone threshold contract, which pays full budget to the agent for every outcome $j$ above some threshold $j^{*}$.
For min-pay rather than min-budget, under MLR we get a monotone contract with a single positive payment, which is optimal (see [14, Lemma 7]). Finally, we note that finding cost-robust min-budget contracts with two levels of pay (“all-or-nothing” [32]) is computationally hard in the general case, as the reduction by [32, Theorem 3] also applies in the cost-robust case (see Section D.8).